The average monthly income of employees in a certain company is P28,000 with a

standard deviation of P9,000. Below which income do we find the lowest 5% of the

nincomps?


1
Expert's answer
2022-03-21T00:29:43-0400

"P(Z<z)=0.05 \to z=-1.645. \;\;\frac{x-28000}{9000}=-1.645 \to x=28000-1.645*9000=13195."

"\frac{}{}"

"\frac{}{}"